# LauraTSE: Target Speaker Extraction using Auto-Regressive Decoder-Only Language Models

### v1

@@ III-C Evaluation Metrics
WavLM input 3.507 3.951 3.137 3.220 0.792 0.447 0.86 0.633
Nemo Conformer 3.621 4.101 3.350 4.243 0.734 1.567 0.929 0.765
III-D Baseline models
We compare LauraTSE with several recent baselines. First, we include Spex+ [3], a discriminative TSE model trained on Libri2Mix. We also compare it with the BSRNN model from WeSep [24] using the results provided in[12]. Additionally, we include AnyEnhance [12], a multi-task model based on masked generative modeling, and we compare with the TSE results.
IV Results and Discussions
Table I presents the overall results on the Libri2Mix Clean test set. LauraTSE achieves comparable speech quality to AnyEnhance, while outperforming it in both speaker similarity and semantic similarity. Notably, LauraTSE is trained on only 460 hours of data, whereas AnyEnhance is trained on 5000 hours, raising the question of whether multi-task learning is always superior to task-specific models for a particular objective. Compared with TSELM [16], which uses discrete tokens from WavLM, our model leverages neural audio codec representations that better preserve speaker identity, addressing the issue of low speaker similarity. Additionally, LauraTSE outperforms discriminative baselines such as Spex+ [3] and WeSep [24] in terms of speech quality, while maintaining competitive performance in semantic similarity and speaker consistency. Table II shows the results of ablation studies using models trained on Libri2Mix. ”Base” refers to the proposed lauraTSE model. ”Nq-” refers to the output number of layers of the AR model. Changing $n_{q}$ from 1 to 3 results in minimal performance differences, suggesting that even a small number of coarse layers may provide sufficient information for the AR model.
Auto-regressive decoder-only model like LauraGPT [18] conducts tasks like SE where the output is strictly aligned with the input length. Therefore, following their approach, we have formatted the decoder-only input to be $[<\text{bos}>,r,m,<\text{tse}>]$. Unlike the original method, which only produces the clean speech, this approach concatenates the reference and the mixture speech into a single input sequence, expecting the model to generate an output containing both the reference speech and the enhanced speech. This method is referred to as the ”Ref output.” During inference, we retain only the portion of the output corresponding to the mixture, constrained by the length of the reference speech. This approach yields results similar to the original one but has led to some undesirable cases where the output length is zero. In contrast, our original method demonstrates that the output sequence can be just the clean speech and does not need to align with the continuous input condition sequence.
@@ IV Results and Discussions
Following the approach in SpeechX [17], we conduct experiments where the input sequences consist of discrete token embeddings rather than continuous log-mel spectrograms, referred to as ”Discrete IO.” For both the input reference speech and the mixture speech, instead of using continuous log-mel spectrograms, we utilize the first two layers of audio codec. We then use two learnable embedding matrices to embed the discrete tokens and summarize these embeddings into a single embedding. This single embedding is then fed into the AR model. We observe that these discretized representations perform worse than the continuous approach. One possible reason for this could be that the neural audio codec representations may not be optimal for usage with our small-scale AR model. Additionally, since the audio codec is typically trained on clean speech, it might miss crucial information from the mixture. A potential direction for future work could involve developing audio codecs that can effectively handle mixture speech.
@@
We also utilize WavLM features instead of the neural audio codec as the embedding features. We use the output from the 6th hidden layer of WavLM as input features for both the reference and the mixture. Additionally, we apply the concatenation technique used in TSELM [16] for mixture representations. Following the approach in SELM [19], the output of the AR model is the Kmeans discrete representations of the target speech. After obtaining the target discrete embeddings, we use a conformer detokenizer to reconstruct the continuous embeddings, as done in SELM [19]. Similar as [16], this approach results in poor speaker similarity, likely due to the discretization process that loses speaker information. In addition, we observe that the semantic similarity is also low, raising the question of how to effectively integrate self-supervised models within AR models.
It has been shown that using a pretrained ASR conformer is beneficial for the Speaker Verification task [35]. We apply the pretrained Nemo Conformer from the Nemo Toolkit [36] to the input mixture and reference (denoted as ”Nemo Conformer”). However, this approach does not outperform the original one and yields poor dWER results. Some problematic cases lead to infinite looping. One possible reason could be the batch normalization problem as stated in [18], while another might be the limited amount of training data. Further research is needed to address these issues.
@@ V Conclusion
Acknowledgment
This research is funded by DKU foundation project ”Emerging AI Technologies for Natural Language Processing”. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.

### v3

@@ IV-B Data Scalability of LauraTSE
WavLM input 3.507 3.951 3.137 3.220 0.792 0.447 0.860 0.633
IV-C What role does the Encoder-only LM play?
Our ultimate goal is to perform TSE using only a decoder-only LM, without relying on an additional encoder-only LM. However, it is difficult for the single-layer neural audio codec to preserve fine-grained speech details as well as multi-layer codecs. Therefore, we need to add the encoder-only LM to refine the coarse-grained output from the decoder-only LM into high-resolution, continuous embeddings. This section studies the effect of the encoder-only LM.
IV-C1 Joint vs. Separate Training
Firstly, we are interested in whether decoder-only LM and encoder-only LM needs to be trained jointly as their tasks are different. We conduct the following experiments: Decoder-Encoder-joint and Decoder-Encoder-split. Details can be found at Figure 5. Decoder-Encoder-joint is the original LauraTSE model where the encoder-only LM and the decoder-only LM are trained jointly through a Straight-Through Estimator (STE) module [41] which ensures the gradient flow for softmax. For Decoder-Encoder-split, the decoder-only LM and the encoder-only LM are trained separately. The results can be found at Table II. We find that training decoder-only LM and encoder-only LM separately achieves slightly superior performance compared with training them jointly, suggesting it not necessary to train them together.
IV-C2 Input composition for Encoder-only LM
The current input to the encoder-only LM contains both the mixture and the reference. To fully study the role of the encoder-only LM, we need to study which input speech information is more important. Therefore, we compose three systems: Encoder-All, Encoder-Mix, and Encoder-Ref. Encoder-All is the original encoder that takes both the mixture and the reference. For Encoder-Mix, only the mixture embedding is fed to the encoder, while Encoder-Ref only takes the reference embedding. Details can be found at Figure 4.
The results are shown at Table III. We can see that the Encoder-Mix achieves similar performance as Encoder-All, while Encoder-Ref achieves much worse performance. This result indicates that the mixture embedding is crucial for the encoder-only LM, and the reference embedding might not be necessary here.
IV-C3 Does Encoder-only LM help the TSE?
Given that the mixture representation for the encoder-only LM is crucial for the final results, we assume that the encoder-only LM not only generates the fine-grained representation from the coarse output of the decoder-only LM, but more importantly, it also aids the TSE task. To further validate this, we compare:
  • Target-n: Clean target speech reconstructed using only the first $n$ layers of the codec. Same as the original LauraTSE, we set $n$ as 2.
  • No-Encoder: Speech reconstructed from the first $n$ layers output codec from the decoder-only LM only.
  • Decoder-Encoder-split: The encoder-only LM is applied after the prediction from the decoder-only LM.
As shown in Table II, No-Encoder underperforms compared to Target-n, indicating that the decoder-only LM alone is insufficient for this task. Adding back the encoder-only LM leads to a significant improvement (Decoder-Encoder-split), reducing the dWER from 0.579 to 0.232.
These findings suggest that the encoder-only LM is currently essential—not only for reconstructing fine-grained speech features but also for assisting with the TSE task. Our future works will focus on developing techniques that allow the decoder-only LM to perform TSE effectively without relying on an additional encoder.
IV-D Other Ablation Studies
Table IV shows the results of other ablation studies we conduct. Base refers to the proposed LauraTSE model. $n$- refers to the output number of layers of the AR decoder-only LM. Changing $n$ from 1 to 3 results in minimal performance differences, suggesting that even a small number of coarse layers may provide sufficient information for TSE.
@@ IV-D Other Ablation Studies
AR decoder-only LMs like LauraGPT [18] conduct tasks like SE where the output is strictly aligned with the input length. To study if the output length has to be the same as the input, we have formatted the decoder-only input to be $\textit{bos},E_{r},E_{m},\textit{tse}]$. Unlike the original method, which only produces the clean speech, this approach concatenates the reference and the mixture speech into a single input sequence, expecting the model to generate an output containing both the reference speech and the enhanced speech. This method is referred to as the Ref output. During inference, we retain only the portion of the output corresponding to the mixture, constrained by the length of the reference speech. This approach yields results similar to the original one. Therefore, we conclude that the output sequence can be just the clean speech and does not need to align with the continuous input condition sequence. Following the approach in SpeechX [17], we conduct experiments where the input sequences are discrete token embeddings rather than continuous log-mel spectrograms, referred to as Discrete IO. For both the input reference speech and the mixture speech, instead of using continuous log-mel spectrograms, we utilize the first two layers of the RVQ output of the neural audio codec. We then use two learnable embedding matrices to embed the discrete tokens and summarize these embeddings into a single embedding. This single embedding is then fed into the AR model. We observe that these discretized representations perform worse than the continuous approach. One possible reason for this could be that the neural audio codec’s discrete representation may not be optimal for our TSE task with a small-scale decoder-only LM. Additionally, since the audio codec is typically trained on clean speech, it might miss crucial information from the mixture. A potential future work could involve developing audio codecs that can effectively handle mixture speech.
@@
We also utilize WavLM features instead of the neural audio codec as the embedding features. We use the output from the 6th hidden layer of WavLM as input features for both the reference and the mixture. Additionally, we apply the concatenation technique used in TSELM [15] for mixture representations. Following the approach in SELM [19], the output of the AR model is the Kmeans discrete representations of the target speech. After obtaining the target discrete embeddings, we use a conformer detokenizer to reconstruct the continuous embeddings, as done in SELM [19]. Similar as [15], this approach, referred to as WavLM input, results in poor speaker similarity, likely due to the discretization process that loses speaker information.